# On $\mathrm{F}$-spaces of almost-Lebesgue functions

### v2

@@ 2.2. Metrizability of asymptotic $L_{p}$-convergence
=\ = $2^{p-1}\left(\int_{X}|\min(|\lambda_{n}||f_{n}-f|,1)|^{p}\,\mathrm{d}\mu+\|(\lambda_{n}-\lambda)f\|_{\alpha_{p}}^{p}\right)$
@@
Since the function $f_{N}$ belongs to $\Lambda_{p}(\mathrm{X}),$ there exists a measurable set $F_{\delta}$ with $\mu(F_{\delta})<\delta/2$ such that
@@
$<1+2^{p-1}\int_{F_{\delta}^{c}}|g_{N}|^{p}\,\mathrm{d}\mu<\infty$
from which we conclude that $f\in\Lambda_{p}(\mathrm{X})$, as desired. ∎
@@
We show that $L_{0}(\mathrm{X})\subseteq\Lambda_{p}(\mathrm{X})$. Let $f\in L_{0}(\mathrm{X})$ and note that, by definition, $f$ is finite $\mu$-almost everywhere. For each $n\in\mathbb{N}$, let $E_{n}$ be the set $E_{n}=\{x\in X\ |\ |f(x)|^{p}\geq n\}$. It is clear that $E_{n+1}\subseteq E_{n}$ for every natural $n$. We claim that $\mu(E_{n})\to 0$ as $n\to\infty$. Suppose not, then there exists $\varepsilon>0$ such that for every $n\in\mathbb{N}$ there exists $\tilde{k}_{n}\geq n$ so that $\mu(E_{k_{n}})\geq\varepsilon$. Set $k_{1}=\tilde{k}_{1}$. If $\tilde{k}_{2}\geq k_{1}$, set $k_{2}=\tilde{k}_{2}$; otherwise $k_{2}=k_{1}$. Similarly, if $\tilde{k}_{3}\geq k_{2}$, set $k_{3}=\tilde{k}_{3}$; otherwise $k_{3}=k_{2}$. Continuing this process, we find a sequence $(k_{n})\to\infty$ as $n\to\infty$, such that $k_{n+1}\geq k_{n}$ and so $E_{k_{n+1}}\subseteq E_{k_{n}}$ Then
$\lim_{n\to\infty}\mu(E_{k_{n}})=\mu\left(\bigcap_{n\in\mathbb{N}}E_{k_{n}}\right)\geq\varepsilon$
@@ 4. Dominated convergence results
The first result of this section is a Lebesgue dominated convergence theorem for the space $\Lambda_{p}(\mathrm{X})$, being a straightforward adaptation of [4, Theorem 5.6].
@@
Let $(f_{n})\subseteq\Lambda_{1}(\mathrm{X})$ and $f$ be measurable. Assume that $(f_{n})$ converges to $f$ a.e. and that there exists $g\in\Lambda_{1}(\mathrm{X})$ such that $|f_{n}|\leq g$ a.e. for every $n\in\mathbb{N}$. Then $f\in\Lambda_{1}(\mathrm{X})$ and for every $\delta>0$ there exists a measurable set $E_{\delta}$ with $\mu(E_{\delta})<\delta$ such that
@@
The next result is a consequence of the previous proposition, and the first part of its proof is very similar to that of [4, Theorem 7.2].
@@
Since $f_{n}\to f$ and $\sup_{n}|f_{n}|\leq g$ a.e., it follows that $|f|\leq g$ a.e., and hence $f\in\Lambda_{p}(\mathrm{X})$. Moreover, we have $|f_{n}-f|^{p}\leq 2^{p}g^{p}$ a.e., which together with the facts that $\lim|f_{n}-f|^{p}=0$ a.e. and $2^{p}g^{p}\in A_{1}(\mathrm{X})$, implies, by Proposition 4.1, that for every $\delta>0$ there exists $E_{\delta}$ with $\mu(E_{\delta})<\delta$ such that
@@
Next, we provide a direct proof that convergence almost in $L_{p}$ implies convergence in the $\mathrm{F}$-norm $\|\cdot\|_{\alpha_{p}}$, even though this result can also be easily derived from [2, Proposition 2.9] and Proposition 2.1. Given $\varepsilon>0$, there exist a measurable set $E_{\varepsilon}$, with $\mu(E_{\varepsilon})<\varepsilon/2$, and $N_{\varepsilon}\in\mathbb{N}$ such that
@@
$\|f_{n}-f\|_{\alpha_{p}}^{p}$ $=\int_{X}|\min(|f_{n}-f|,1)|^{p}\,\mathrm{d}\mu$
$=\int_{E_{\varepsilon}}|\min(|f_{n}-f|,1)|^{p}\,\mathrm{d}\mu+\int_{E_{\varepsilon}^{c}}|\min(|f_{n}-f|,1)|^{p}\,\mathrm{d}\mu$
$\leq\mu(E_{\varepsilon})+\int_{E_{\varepsilon}^{c}}|f_{n}-f|^{p}\,\mathrm{d}\mu<\varepsilon$
which finishes the proof. ∎
The previous result also holds if convergence almost everywhere is replaced by convergence in measure, serving as a counterpart to [4, Theorem 7.8] in the present setting.
@@ 5. Vitali convergence theorems
To establish the converse, we show that $(f_{n})$ is Cauchy in $L_{p}(\mathrm{X})$. Since convergence almost in $L_{p}$ implies $\alpha_{p}$-convergence, it suffices to assume the latter to establish the result. Suppose that $(f_{n})$ is uniformly $p$-integrable and let $\varepsilon>0$ and $\delta_{\varepsilon}$ be such that $\sup_{n}\int_{F}|f_{n}|^{p}\,\mathrm{d}\mu<\varepsilon^{p}$ whenever $\mu(F)<\delta_{\varepsilon}$. Assume that $(f_{n})$ $\alpha_{p}$-converges to $f$. Then, $(f_{n})$ is $\alpha_{p}$-Cauchy and hence there exists a sequence of measurable sets $(B_{n})$ with $\mu(B_{n}^{c})\to 0$ such that
@@
First, assume that $(f_{n})$ $\alpha_{p}$-converges to $f$ (hence $f\in\Lambda_{p}(\mathrm{X})$) and let $\varepsilon>0$. Since $\alpha_{p}$-convergence implies convergence in measure [2, 1], we put our attention on the second condition. Let $N=N(\varepsilon)\in\mathbb{N}$ be such that $\|f_{n}-f\|_{\alpha_{p}}<\varepsilon/2$ whenever $n>N$. Since $f\in\Lambda_{p}(\mathrm{X})$, there exists a measurable set $F_{\varepsilon}$ with $\mu(F_{\varepsilon})<\varepsilon$ so that $f\chi_{F_{\varepsilon}^{c}}\in L_{p}(\mathrm{X})$. Then, $\|f\chi_{F_{\varepsilon}^{c}\cap G_{\varepsilon}^{c}}\|_{p}<\varepsilon/2$ for some subset $G_{\varepsilon}\subseteq\mathrm{X}$ of finite measure. Therefore, for $n>N$,
@@
$\leq\|f_{n}-f\|_{\alpha_{p}}+\|f\chi_{F_{\varepsilon}^{c}\cap G_{\varepsilon}^{c}}\|_{p}<\varepsilon.$
For each $n=1,\ldots,N$, since $f_{n}\in\Lambda_{p}(\mathrm{X})$, there exists $H_{\varepsilon,n}$ with $\mu(H_{\varepsilon,n})<\varepsilon$ such that $f_{n}\chi_{H_{\varepsilon,n}}\in L_{p}(\mathrm{X})$, and so $\|f_{n}\chi_{H_{\varepsilon,n}^{c}\cap I_{\varepsilon,n}^{c}}\|_{p}<\varepsilon$ for some $I_{\varepsilon,n}$ with finite measure. Define a measurable set $E_{\varepsilon}$ as
@@
Now, assume that $(f_{n})$ converges to $f$ in measure and that condition $(ii)$ holds. We prove that $(f_{n})$ is $\alpha_{p}$-Cauchy. Let $\varepsilon>0$ and let a measurable set $E_{\varepsilon}$ be such that $\mu(E_{\varepsilon})<\infty$ and $\|f_{n}\chi_{E_{\varepsilon}^{c}}\|_{\alpha_{p}}<\varepsilon/4$ for each $n\in\mathbb{N}$. Note that, for every $n,m\in\mathbb{N}$,
@@
Set $\delta=\varepsilon/2\mu(E_{\varepsilon})^{1/p}$ and let $N_{\varepsilon}\in\mathbb{N}$ be such that $\mu(|f_{n}-f_{m}|\geq\delta)<\varepsilon/2$ whenever $n,m\geq N_{\varepsilon}$. Then, for this range of $n$ and $m$ we estimate
$\|(f_{n}-f_{m})\chi_{E_{\varepsilon}}\|_{\alpha_{p}}^{p}=$ $\int_{E_{\varepsilon}}|\min(f_{n}-f_{m},1)|^{p}\,\mathrm{d}\mu$
=\ = $\int_{E_{\varepsilon}\cap\{|f_{n}-f_{m}|>\delta\}}|\min(f_{n}-f_{m},1)|^{p}\,\mathrm{d}\mu$
$+\int_{E_{\varepsilon}\cap\{|f_{n}-f_{m}|\leq\delta\}}|\min(f_{n}-f_{m},1)|^{p}\,\mathrm{d}\mu$
[…]
@@
$\|f_{n}-f_{m}\|_{\alpha_{p}}<\varepsilon$
@@
If $\mu(\mathrm{X})$ is finite, then condition $(ii)$ of Theorem 5.4 is automatically satisfied by setting $E_{\varepsilon}=\mathrm{X}$ for every $\varepsilon>0$. This, together with Lemma 3.3, establishes the equivalence between asymptotic $L_{p}$-convergence and convergence in measure on finite measure spaces, a result initially proved in [1].
@@ 6. Approximation and separability
The first result of this section establishes the density of simple functions in $\Lambda_{p}(\mathrm{X})$, as a natural extension of [10, Proposition 6.7] to our setting.
@@
Note that $s\in\mathcal{S}\cap\Lambda_{p}(\mathrm{X})$ if and only if for every $\delta>0$ there exists a measurable set $E_{\delta}$ with $\mu(E_{\delta})<\delta$ such that the set $\{s\neq 0\}\cap E_{\delta}^{c}$ has finite measure.
The next result establishes the density of $L_{p}(\mathrm{X})$ in $\Lambda_{p}(\mathrm{X})$.
@@
Let $f\in\Lambda_{p}(\mathrm{X})$ and $\varepsilon>0$. There exists a measurable set $E_{\varepsilon}$ with $\mu(E_{\varepsilon})<\varepsilon^{p}$ such that $g_{\varepsilon}=f\chi_{E_{\varepsilon}^{c}}$ belongs to $L_{p}(\mathrm{X})$. We have
$\|f-g_{\varepsilon}\|_{\alpha_{p}}^{p}$ $=\int_{\mathrm{X}}|\min(|f-g_{\varepsilon}|)|^{p}\,\mathrm{d}\mu$
$=\int_{E_{\varepsilon}^{c}}|\min(|f-f\chi_{E_{\varepsilon}^{c}}|,1)|^{p}\,\mathrm{d}\mu+\int_{E_{\varepsilon}}|\min(|f-f\chi_{E_{\varepsilon}^{c}}|,1)|^{p}\,\mathrm{d}\mu.$

### v3

@@ 2.4. Topology of (local) convergence in measure
=\ = $2^{p-1}\left(\int_{X}\min(|\lambda_{n}||f_{n}-f|,1)^{p}\,\mathrm{d}\mu\right)$
$+2^{p-1}\left(\int_{X}\min(|\lambda_{n}-\lambda||f|,1)^{p}\,\mathrm{d}\mu\right)$
@@
Since the function $g_{N}$ belongs to $\Lambda_{p}(\mathrm{X}),$ there exists a measurable set $F_{\delta}$ with $\mu(F_{\delta})<\delta/2$ such that
@@
$<1+2^{p-1}\int_{F_{\delta}^{c}}|g_{N}|^{p}\,\mathrm{d}\mu$
<$\infty$< ∞
from which it follows that $f\in\Lambda_{p}(\mathrm{X})$, as desired. ∎
@@
We show that $L_{0}(\mathrm{X})\subseteq\Lambda_{p}(\mathrm{X})$. Let $f\in L_{0}(\mathrm{X})$ and note that, by definition, $f$ is finite $\mu$-almost everywhere. For each $n\in\mathbb{N}$, let $E_{n}$ be the set $E_{n}=\{x\in X\ |\ |f(x)|^{p}>n\}$. It is clear that $E_{m}\subseteq E_{n}$ for all $n,m\in\mathbb{N}$ such that $m\geq n$. We claim that $\mu(E_{n})\to 0$ as $n\to\infty$. Suppose not, then there exists $\varepsilon>0$ such that for every $n\in\mathbb{N}$ there exists $\tilde{k}_{n}\geq n$ so that $\mu(E_{\tilde{k}_{n}})\geq\varepsilon$. Set $k_{1}=\tilde{k}_{1}$. If $\tilde{k}_{2}\geq k_{1}$, set $k_{2}=\tilde{k}_{2}$; otherwise $k_{2}=k_{1}$. Similarly, if $\tilde{k}_{3}\geq k_{2}$, set $k_{3}=\tilde{k}_{3}$; otherwise $k_{3}=k_{2}$. Continuing this process, we find a sequence $(k_{n})\to\infty$ as $n\to\infty$, such that $k_{n+1}\geq k_{n}$ and so $E_{k_{n+1}}\subseteq E_{k_{n}}$. Then
$\mu\left(\bigcap_{n\in\mathbb{N}}E_{k_{n}}\right)=\lim_{n\to\infty}\mu(E_{k_{n}})\geq\varepsilon$
@@ 4. Dominated convergence results
The first result of this section is a Lebesgue dominated convergence theorem for the space $\Lambda_{p}(\mathrm{X})$, being a straightforward adaptation of [3, Theorem 5.6].
@@
Let $(f_{n})\subseteq\Lambda_{1}(\mathrm{X})$ and $f$ be measurable. Assume that $(f_{n})$ converges to $f$ almost everywhere and that there exists $g\in\Lambda_{1}(\mathrm{X})$ such that $\sup_{n}|f_{n}|\leq g$ almost everywhere. Then $f\in\Lambda_{1}(\mathrm{X})$ and for every $\delta>0$ there exists a measurable set $E_{\delta}$ with $\mu(E_{\delta})<\delta$ such that
@@
Proof.
Let $\delta>0$ and $E_{\delta}$ be a measurable set with $\mu(E_{\delta})<\delta$ such that $g\chi_{E_{\delta}^{c}}\in L_{p}(\mathrm{X})$. Since $|f_{n}|\leq g$ and $(f_{n})$ converges to $f$ almost everywhere, then $|f|\leq g$ and $g+f_{n}$, $g-f_{n}$ are nonnegative almost everywhere. It follows that $f\in\Lambda_{1}(\mathrm{X})$ and by Fatou’s lemma we have:
$\int_{E_{\delta}^{c}}g\,\mathrm{d}\mu+\int_{E_{\delta}^{c}}f\,\mathrm{d}\mu\leq\int_{E_{\delta}^{c}}g\,\mathrm{d}\mu+\liminf\limits_{n\to\infty}\int_{E_{\delta}^{c}}f_{n}\,\mathrm{d}\mu$
and, similarly,
$\int_{E_{\delta}^{c}}g\,\mathrm{d}\mu-\int_{E_{\delta}^{c}}f\,\mathrm{d}\mu$ $\leq\int_{E_{\delta}^{c}}g\,\mathrm{d}\mu-\limsup\limits_{n\to\infty}\int_{E_{\delta}^{c}}f_{n}\,\mathrm{d}\mu$
whence
$\limsup\limits_{n\to\infty}\int_{E_{\delta}^{c}}f_{n}\,\mathrm{d}\mu\leq\int_{E_{\delta}^{c}}f\,\mathrm{d}\mu\leq\liminf\limits_{n\to\infty}\int_{E_{\delta}^{c}}f_{n}\,\mathrm{d}\mu$
establishing the result. ∎
The next result is a consequence of the previous proposition, and the first part of its proof is very similar to that of [3, Theorem 7.2].
@@
Since $f_{n}\to f$ and $\sup_{n}|f_{n}|\leq g$ a.e., it follows that $|f|\leq g$ a.e., and hence $f\in\Lambda_{p}(\mathrm{X})$. Moreover, we have $|f_{n}-f|^{p}\leq 2^{p}g^{p}$ a.e., which together with the facts that $\lim|f_{n}-f|^{p}=0$ a.e. and $2^{p}g^{p}\in\Lambda_{1}(\mathrm{X})$, implies, by Proposition 4.1, that for every $\delta>0$ there exists $E_{\delta}$ with $\mu(E_{\delta})<\delta$ such that
@@
Next, we provide a direct proof that convergence almost in $L_{p}$ implies convergence in the $\mathrm{F}$-norm $\|\cdot\|_{\alpha_{p}}$; alternatively one could apply [2, Proposition 2.9] and Proposition 2.1. Given $\varepsilon>0$, there exist a measurable set $E_{\varepsilon}$, with $\mu(E_{\varepsilon})<\varepsilon/2$, and $N_{\varepsilon}\in\mathbb{N}$ such that
@@
$\|f_{n}-f\|_{\alpha_{p}}^{p}$ $=\int_{X}\min(|f_{n}-f|,1)^{p}\,\mathrm{d}\mu$
$=\int_{E_{\varepsilon}}\min(|f_{n}-f|,1)^{p}\,\mathrm{d}\mu+\int_{E_{\varepsilon}^{c}}\min(|f_{n}-f|,1)^{p}\,\mathrm{d}\mu$
$\leq\mu(E_{\varepsilon})+\int_{E_{\varepsilon}^{c}}|f_{n}-f|^{p}\,\mathrm{d}\mu$
$<\varepsilon$
as required. ∎
The previous result also holds if convergence almost everywhere is replaced by convergence in measure, serving as a counterpart to [3, Theorem 7.8] in the present setting.
@@
Proof.
Suppose that $(f_{n})$ does not $\alpha_{p}$-converge to $f$. Then, there exist $\varepsilon>0$ and a subsequence $(g_{n})$ of $(f_{n})$ such that $\|g_{n}-f\|_{\alpha_{p}}\geq\varepsilon$. By hypothesis, $(g_{n})$ converges to $f$ in measure, and hence there exists a subsequence $(h_{n})$ of $(g_{n})$ converging to $f$ almost everywhere. Given that $(h_{n})$ is dominated by $g$, it follows by Proposition 4.2 that $(h_{n})$ converges to $f$ with respect to $\|\cdot\|_{\alpha_{p}}$, a contradiction. Therefore, $(f_{n})$ converges to $f$ in the $\mathrm{F}$-norm $\|\cdot\|_{\alpha_{p}}$, which implies that $f\in A_{p}(\mathrm{X})$. ∎
@@ 5. Vitali convergence theorems
To establish the converse, we show that $(f_{n})$ is Cauchy in $L_{p}$. Since convergence almost in $L_{p}$ implies $\alpha_{p}$-convergence, it suffices to assume the latter to establish the result. Suppose that $(f_{n})$ is uniformly $p$-integrable and let $\varepsilon>0$ and $\delta_{\varepsilon}$ be such that $\sup_{n}\int_{F}|f_{n}|^{p}\,\mathrm{d}\mu<\varepsilon^{p}$ whenever $\mu(F)<\delta_{\varepsilon}$. Assume that $(f_{n})$ $\alpha_{p}$-converges to $f$. Then, $(f_{n})$ is $\alpha_{p}$-Cauchy and hence there exists a sequence of measurable sets $(B_{n})$ with $\mu(B_{n}^{c})\to 0$ such that
@@
First, assume that $(f_{n})$ $\alpha_{p}$-converges to $f$ (hence $f\in\Lambda_{p}(\mathrm{X})$) and let $\varepsilon>0$. Since $\alpha_{p}$-convergence implies (local) convergence in measure, we put our attention on the second condition. Let $N=N(\varepsilon)\in\mathbb{N}$ be such that $\|f_{n}-f\|_{\alpha_{p}}<\varepsilon/2$ whenever $n>N$. Since $f\in\Lambda_{p}(\mathrm{X})$, there exists a measurable set $F_{\varepsilon}$ with $\mu(F_{\varepsilon})<\varepsilon$ so that $f\chi_{F_{\varepsilon}^{c}}\in L_{p}(\mathrm{X})$. Then, $\|f\chi_{F_{\varepsilon}^{c}\cap G_{\varepsilon}^{c}}\|_{p}<\varepsilon/2$ for some subset $G_{\varepsilon}\subseteq\mathrm{X}$ of finite measure. Therefore, for $n>N$,
@@
$\leq\|f_{n}-f\|_{\alpha_{p}}+\|f\chi_{F_{\varepsilon}^{c}\cap G_{\varepsilon}^{c}}\|_{p}$
$<\varepsilon\,.$
For each $n=1,\ldots,N$, since $f_{n}\in\Lambda_{p}(\mathrm{X})$, there exists $H_{\varepsilon,n}$ with $\mu(H_{\varepsilon,n})<\varepsilon$ such that $f_{n}\chi_{H_{\varepsilon,n}^{c}}\in L_{p}(\mathrm{X})$, and so $\|f_{n}\chi_{H_{\varepsilon,n}^{c}\cap I_{\varepsilon,n}^{c}}\|_{p}<\varepsilon$ for some $I_{\varepsilon,n}$ of finite measure. Define a measurable set $E_{\varepsilon}$ as
@@
Now, assume that $(f_{n})$ converges to $f$ locally in measure and that condition $(ii)$ holds. We prove that $(f_{n})$ is $\alpha_{p}$-Cauchy. Let $\varepsilon>0$ and let $E_{\varepsilon}\in\Sigma$ be such that $\mu(E_{\varepsilon})<\infty$ and $\|f_{n}\chi_{E_{\varepsilon}^{c}}\|_{\alpha_{p}}<\varepsilon/4$ for each $n\in\mathbb{N}$. Note that, for every $n,m\in\mathbb{N}$,
@@
Moreover, since $\mu(E_{\varepsilon})<\infty$, the local convergence in measure of $(f_{n})$ towards $f$ yields the existence of $N_{\varepsilon}\in\mathbb{N}$ so that $\|f_{n}-f_{m}\|_{\alpha,E_{\varepsilon}}<(\varepsilon/2)^{p}$ for all $n,m\geq N_{\varepsilon}$.
@@
$\|f_{n}-f_{m}\|_{\alpha_{p}}$ $<\|(f_{n}-f_{m})\chi_{E_{\varepsilon}}\|_{\alpha_{p}}+\varepsilon/2$
$\leq\|f_{n}-f_{m}\|_{\alpha,E_{\varepsilon}}^{1/p}+\varepsilon/2$
$<\varepsilon$
@@
If $\mu(\mathrm{X})$ is finite, then condition $(ii)$ of Theorem 5.4 is automatically satisfied by setting $E_{\varepsilon}=\mathrm{X}$ for every $\varepsilon>0$, thereby recovering the equivalence between asymptotic $L_{p}$-convergence and (local) convergence in measure on finite measure spaces.
@@ 6. Approximation and separability
The first result of this section establishes the density of simple functions in $\Lambda_{p}(\mathrm{X})$, as a natural extension of [9, Proposition 6.7] to our setting.
@@
Proof.
Let $f\in\Lambda_{p}(\mathrm{X})$. By applying [3, Lemma 2.11] to the positive and negative parts of $f$, we construct a sequence of simple functions $(s_{n})$ that converges to $f$ pointwise and $\sup_{n}|s_{n}|\leq|f|$. Clearly, $(s_{n})\subseteq\Lambda_{p}(\mathrm{X})$. Additionally, $|s_{n}-f|^{p}\leq 2^{p}|f|^{p}\in\Lambda_{1}(\mathrm{X})$ and $|s_{n}-f|^{p}\to 0$ pointwise, and hence, by Proposition 4.1, we conclude that $(s_{n})$ converges to $f$ almost in $L_{p}$. The result follows. ∎
Note that $s\in\mathcal{S}\cap\Lambda_{p}(\mathrm{X})$ if, and only if, for every $\delta>0$ there exists a measurable set $E_{\delta}$ with $\mu(E_{\delta})<\delta$ such that the set $\{s\neq 0\}\cap E_{\delta}^{c}$ has finite measure.
The next result establishes the density of $L_{p}(\mathrm{X})$ in $\Lambda_{p}(\mathrm{X})$. Together with Theorem 1.1, this implies that on finite measure spaces, $L_{p}(\mathrm{X})$ is dense in $L_{0}(\mathrm{X})$ with respect to the topology of (local) convergence in measure.
@@
Let $f\in\Lambda_{p}(\mathrm{X})$ and $\varepsilon>0$. There exists a measurable set $E_{\varepsilon}$ with $\mu(E_{\varepsilon})<\varepsilon^{p}$ such that $g_{\varepsilon}=f\chi_{E_{\varepsilon}^{c}}$ belongs to $L_{p}(\mathrm{X})$. We have:
$\|f-g_{\varepsilon}\|_{\alpha_{p}}^{p}$ $=\int_{\mathrm{X}}\min(|f-g_{\varepsilon}|,1)^{p}\,\mathrm{d}\mu$
$=\int_{E_{\varepsilon}^{c}}\min(|f-f\chi_{E_{\varepsilon}^{c}}|,1)^{p}\,\mathrm{d}\mu+\int_{E_{\varepsilon}}\min(|f-f\chi_{E_{\varepsilon}^{c}}|,1)^{p}\,\mathrm{d}\mu\,.$